OpenStudy (anonymous):

solve for x 3/2(x - 5)

OpenStudy (idku):

this is supposed to be equal to something?

OpenStudy (anonymous):

sorry yes = (-6)

OpenStudy (idku):

\(\dfrac{3}{2(x-5)}=-6\)

OpenStudy (idku):

this /

OpenStudy (idku):

Or is it \(\dfrac{3}{2}(x-5)=-6\) ?

OpenStudy (anonymous):

its 3/2 is a fraction not dividing

OpenStudy (anonymous):

2nd one

OpenStudy (idku):

so it is the second thing i said... nice. I would do this (1) Multiply both sides by 2 (2) Divide both sides by 3 (3) Add 5 to both sides

OpenStudy (anonymous):

ok ill try that thanks

OpenStudy (anonymous):